### v2

# Not all Layers of LLMs are Necessary during Inference

Abstract
The inference phase of Large Language Models (LLMs) is very expensive. An ideal inference stage of LLMs could utilize fewer computational resources while still maintaining its capabilities (e.g., generalization and in-context learning ability). In this paper, we try to answer the question, “During LLM inference, can we use shallow layers for easy instances; and deep layers for hard ones?” To answer this question, we first indicate that Not all Layers are Necessary during Inference by statistically analyzing the activated layers across tasks. Then, we propose a simple algorithm named AdaInfer to determine the inference termination moment based on the input instance adaptively. More importantly, AdaInfer does not alter LLM parameters and maintains generalizability across tasks. Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer saves an average of 14.8% of computational resources, even up to 50% on sentiment tasks, while maintaining comparable performance. Additionally, this method is orthogonal to other model acceleration techniques, potentially boosting inference efficiency further.
Not all Layers of LLMs are Necessary during Inference
Siqi Fan2, Xin Jiang1, Xiang Li1, Xuying Meng3, Peng Han2, Shuo Shang2*, Aixin Sun4, Yequan Wang1*, Zhongyuan Wang1 1Beijing Academy of Artificial Intelligence, Beijing, China 2University of Electronic Science and Technology of China, Chengdu, China 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 4School of Computer Science and Engineering, Nanyang Technological University, Singapore
11footnotetext: Corresponding authors.
1 Introduction
LLMs have demonstrated impressive performance on various downstream tasks (e.g., text generation, question & answering, and sentiment analysis) using various evaluation protocols such as zero-shot, few-shot, and fine-tuning Todd et al. (2024); Chan et al. (2022); Kossen et al. (2023); Wang et al. (2023, 2022). Notably, In-context learning ability allows LLMs to adapt to tasks using input-output examples without parameter updates Kossen et al. (2023); Todd et al. (2024). However, their inference phases are very expensive Pope et al. (2023); Liu et al. (2023). For example, the inference time complexity for typical large models with Transformer structure is $LSd(d+S)$ per single inference, where $d$, $S$, and $L$ represent the hidden size, sequence length, and layer number, respectively. An ideal inference LLM should utilize fewer computational resources while still maintaining its capabilities in generalization and in-context learning ability Liu et al. (2023). The popular methods for achieving efficient inference in LLMs include model pruning Liu et al. (2018) and sparse models LeCun et al. (1989). However, altering LLM parameters may risk compromising its generalization ability, which is challenging to detect. Meanwhile, different LLM designs pose compatibility challenges with other acceleration methods.
In this paper, we consider dynamically reducing the number of activated neurons as an approach to accelerate LLM inference. Inspired by the human thinking process Salthouse (1996); Deary et al. (2001), where quick answers are often provided for simple questions while more time is spent on thoughtful reasoning for complex ones, e.g., knowledge-related questions. Previous studies Teerapittayanon et al. (2016); Huang et al. (2017) show that “Easy” tasks activate at shallower layers while “hard” ones at deeper layers. Additionally, growth strategy Li et al. (2023) is proposed to lower the training cost of LLMs by adding parameters in stages. It inspires us that reducing the computing parameters during inference may be an effective way besides existing typical accumulation methods. Statistical LLMs results on various tasks (see Section 3.2 for detail) show that reducing parameters is feasible during LLM inference.
Therefore, a natural approach to achieve LLM efficient inference is to decide when to stop the inference process based on the input instance adaptively. For instance, allocating fewer computational resources for processing “simple” samples to enhance operational efficiency. Furthermore, exploring adaptive inference may bridge LLMs with the brain’s information processing Hubel and Wiesel (1962); Murata et al. (2000), aiding in the analysis of activated network modules during sample processing Han et al. (2021) and identifying crucial input components affecting the final prediction.
Specifically, we present AdaInfer, a simple but effective algorithm for instance-aware adaptive inference. The core of AdaInfer lies in data-driven decision-making. Generally, there are two approaches to getting decision-making signals: (1) updating LLM parameters requires training, involves high costs, and might decrease the model’s generalizability Gu et al. (2024), and (2) keeping parameters unchanged, a more desirable and cost-effective approach that preserves the model’s innate ability Yao et al. (2023). In this work, we adopt an early stop strategy, optimizing efficiency without altering the model’s parameters. In particular, we begin by performing statistical analysis on LLM for each block feature (e.g., logits, hidden state, mlp, and attention activation value). Subsequently, we choose logits to construct features and employ classical statistical classifiers (i.e., SVM and CRF) to facilitate the early exit strategy (Section 4).
Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer can save an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is orthogonal to other model acceleration techniques, offering the potential for further enhancing inference efficiency (Section 5).
2 Related Work
A straightforward approach to achieve adaptive inference involves dynamic neural networks Han et al. (2021); Huang et al. (2017); Bolukbasi et al. (2017). Consequently, networks with dynamic structures can be classified into two types: dynamic depth (number of network layers) and width (number of channels, parallel subnetworks etc.).
Dynamic depth.
Dynamic depth involves two methods: Early Exit (EE) and Skip layer. EE first appeared in CNN/DNN networks for visual tasks Bolukbasi et al. (2017); Huang et al. (2017); Teerapittayanon et al. (2016). Subsequently, it was utilized in accelerating the inference of encoder-only architectures in BERT by Li et al. (2020); Liu et al. (2020); Li et al. (2021); Kong et al. (2022). Recently, Schuster et al. (2022); Varshney et al. (2023) discuss confidence-based EE for LM adaptive inference. Our proposed AdaInfer closely aligns with EE concept. Specifically, we apply EE to mainstream decoder-only LLMs, which adhere to the scaling law but suffer from high inference costs due to their large parameter count. Meanwhile, skip-layer dynamically omits the execution of middle layers (or modules) for any input token, facilitated by a gate function Wang et al. (2018) or a binary router Zeng et al. (2023) and layer pruning Kim et al. (2024); Yang et al. (2024); Song et al. (2024). The main difference between our method and theirs is that we achieve instance-wise inference without altering the model parameters, which is crucial for current LLMs. To the best of our knowledge, this is the first attempt to discover that each block’s logits are crucial elements for EE classifiers in LLMs, and we incorporate it as a fundamental design choice in AdaInfer.
Dynamic width.
Dynamic width controls the number of neurons on the network width for efficient inference, such as reducing the number of CNN channels Hua et al. (2019); Hoefler et al. (2021) and establishing multiple parallel structures for “experts” in MoE Fedus et al. (2022); Zhou et al. (2022); Artetxe et al. (2021), dynamically weighting and predicting the output results. Recently, Ma et al. (2023); Addanki et al. (2023); Xia et al. (2023) slimming the network width by pruning attention heads and the output neuron in Query, Key or Value. Other model accelerate methods such as quantization Xiao et al. (2023); Frantar et al. (2022), sparsity Liu et al. (2023) are orthogonal areas and usually excel in different settings.
3 Efficiency Analysis of LLM Inference
This section aims to prove that Not all Layers are Necessary during Inference by analyzing the number of activated layers across various tasks. We first briefly review LLM’s critical components. Then, we present our statistical observations and insights.
3.1 Preliminary: LLM Building Blocks
Modern LLMs, rooted in the Transformer architecture Vaswani et al. (2017), are trained with different unsupervised training objectives. For instance, mainstream LLMs (e.g., GPT, Llama series) are pre-trained with a full language modeling objective with a decoder-only structure, computing loss on all tokens. The key components of LLMs can be broken down into the following blocks:
Tokenizer and Embedding Layer.
This block tokenizes input text into numerical vectors, enabling effective processing and analysis of textual data.
Decoder Block.
This block processes numerical vectors through self-attention and feedforward neural networks, enabling the model to focus on (attend to) the most relevant input parts.
Classification Layer.
The LM head layer converts decoder logits into a vocabulary-wide probability distribution to facilitate word prediction.
These blocks facilitate LLMs in efficiently handling NLP downstream tasks, with a primary emphasis on decoder blocks within multi-layer Transformers. For typical large Transformer models, inference complexity is linearly related to the number of decoder layers $L$ and is given by $LSd(d+S)$ per single inference. Consequently, to explore the possibility of skipping intermediate layers in LLMs during inference, we do the following statistics.
Figure 1: LLama2-7B model zero/few-shot performance across all decoder layers: solid line for sentiment analysis while dashed line for MMLU tasks.
3.2 Not all Layers are Necessary
Earlier Transformer models typically comprise 6 decoder layers, while current open-source models, such as Llama2-13B Touvron et al. (2023), feature 40 decoder layers. However, during inference, each input instance for different tasks passes through every block layer by layer until the last layer, prompting us to question: “Can we allocate fewer computational resources per input instance instead of the same substantial budget?” To investigate this, we conduct a statistical analysis to examine the correlation between accuracy and the activation of layers across various tasks. The statistical results are depicted in Figure 1.
Observation 1:
Not all Layers of LLMs are Necessary during Inference: Early Stopping works. In sentiment analysis using the Llama2-13B (40 layers) model, the average activated layer count per input is $21$, with a variance of $5.1$. This observation is intuitive. For instance, simpler inputs like “I like Camera A” activate 16 layers, while more complex inputs like “Camera A is better than Camera B in picture quality” activate 24 layers. The latter sentence introduces a comparative sentiment about the “quality” aspect between Camera A and Camera B, which embodies more complex features, suggesting deeper layers for such complex instances.
Observation 2:
Varying Task Difficulties, Different Activation Layers: Stop Simpler Tasks Sooner, Let Complex Ones Go Deeper. Tasks in the LLM activate different layers, with simpler ones usually at shallower layers and more complex ones at deeper layers. This is shown in Figure 1, which demonstrates the performance of a Llama2-7B model across 32 layers in sentiment analysis Socher et al. (2013) and MMLU Hendrycks et al. (2021). For simple tasks like sentiment classification, accuracy matches that of the final layer by the 24th layer. Conversely, for complex tasks like MMLU, accuracy tends to improve with deeper layers.
Insight.
The observations mentioned above are intuitive. It’s worth noting that similar observations have been made by Teerapittayanon et al. (2016); Huang et al. (2017) for visual tasks in convolutional neural networks. Surprisingly, we have also observed this phenomenon at LLM inference time. By exploiting this phenomenon, we can perform instance-aware adaptive inference for LLMs, dynamically adjusting their structure/parameters for different test samples, thereby achieving superior advantages in inference efficiency and adaptability. Moving forward, we will leverage this observation to implement adaptive inference.
(a) A workflow of AdaInfer processing three input instances, involving two for sentiment analysis and one for a knowledge-based question answering task. It shows that the early-exit moment varies across the instances.
(b) After implementing AdaInfer, LLMs can reduce computational costs through adaptive early-exit strategies.
Figure 2: An illustration of AdaInfer’s processing and computational savings.
4 AdaInfer
The workflow of AdaInfer and the computational efficiencies gained through this method are depicted in Figures 2(a) and 2(b), respectively. The key of AdaInfer is how to find the early stop signal while keeping the original abilities of LLMs. AdaInfer dynamically computes the stopping signal by evaluating critical features (i.e., “gap” and “top prob”). This process involves two main components: a Feature Selection module and a Classifier module. At each layer, the Feature Selection crafts a feature vector for the current input instance. Subsequently, the Classifier (often SVM or CRF for their effectiveness) assesses the stopping signal’s strength. A strong enough signal triggers an early process termination, allowing for the bypass of subsequent decoder layers.
4.1 Feature Selection
As we mentioned before, modifying LLM parameters requires training and incurs high costs. More importantly, it may pose a potential risk of compromising the model’s generalization capabilities and detecting these issues can be challenging Gu et al. (2024). Hence, we embrace a more desirable and cost-effective approach that preserves the model’s innate abilities without altering parameters. AdaInfer utilizes specially designed features (e.g., “gap” and “top prob”), leveraging a statistical classifier for evaluation stopping signal.
Problem: The lack of features for decision-making.
LLMs capture coarse-grained features in their initial layers and develop more detailed, fine-grained representations in subsequent, deeper layers, facilitated by repeated application of multi-head attention mechanisms and the use of residual connections. However, there is a lack of universal-level features to demonstrate that shallow-level representation is sufficient for the current task. Furthermore, these features need to be inherently universal to ensure compatibility across various LLMs.
(a) Llama2 on sentiment
(b) Llama2 on MMLU
Figure 3: Statistics of features within LLMs that vary with the forward layer.
Solution: Logits reflect mutation.
To address this, we conducted a visual analysis of diverse features across the layers within each block of LLMs. Our examination focused specifically on:
  • Gap: Measures the current block’s prediction confidence for the next token, defined as $\text{gap}=P(\text{top token})-P(\text{second token})$, where $P$ represents the probability distribution generated by the current block.
  • Top Prob: Indicates $P(\text{top token})$, the probability estimation by the current block for the most likely next token.
  • Cosine Similarity: Calculated to evaluate the similarity between the features of current and previous block, including attention activation value (attn), multi-layer perceptron outputs (mlp), and hidden states.
These analyses are showcased in Figure 3. In this figure, we observe the following trends: (1) For Llama2-13B with 40 layers Touvron et al. (2023) across sentiment and MMLU tasks, the “gap” and “top prob” gradually increase during the inference phase, stabilizing in the deeper layers. (2) The activation of “gap” and “top prob” varies across layers for different tasks. These phenomenons are also evident in the Llama2-7B, OPT-13B Zhang et al. (2022), and GPT-J Wang and Komatsuzaki (2021) (See Appendix C). This demonstrates “gap” and “top prob” can serve as universal features, indicating the stopping signal. Notably, these two values remain consistent across diverse tasks, suggesting a versatile classifier applicable to various tasks. Factor study in subsequent experiments also shows that other features (e.g., Cosine Similarity) exhibit subtle differences across layers.
4.2 Classifier
The Classifier determines if the signal is compelling enough to warrant an early termination of the process. The rule-based approach heavily relies on rules, and the cost of individually constructing domain-specific features is high  Huang et al. (2017); Yang et al. (2020); Wang et al. (2022). Conversely, the plug-and-play nature of the gating function Lin et al. (2017); Bejnordi et al. (2019) provides greater universality. Nonetheless, discrete decision functions, lacking gradient information, often require specialized training methods.
The trend in Figure 3 indicates classical statistical classification methods can address discrete decision-making problems. We can connect block features to decision-making via a statistical classifier. By classifying general features (i.e., “gap” and “top prob”), we simplify decision-making into binary classification, enabling an early exit strategy. If the classifier considers the current layer’s features stoppable, subsequent layers’ computations can be discarded; otherwise, continue to the final layer. This process is also illustrated in Figure 2(a).
4.3 Classifier Objective
Here we detail the training process of the classifier through their objectives, respectively. Given one instance, we calculate the feature vector $x_{d}$ using the feature selection module. This feature vector serves as the input for the classifier module. If the current layer’s output $\hat{y}$ provides the correct answer $y$, the associated label $y_{c}$ is a positive example; otherwise, it’s a negative example.
$y_{c}=\begin{cases}1&\text{if }\hat{y}=y,\\ 0&\text{otherwise}.\end{cases}$ (1)
Thus, for an $L-$layer LLM, each input instance $x$ yields $L$ pairs of $<x^{d},y_{c}>$. The details of creating training data for classifier are in Appendix B. We consider two types of classifiers, Support Vector Machines (SVM) Hearst et al. (1998) and Conditional Random Fields (CRF) Lafferty et al. (2001). The first one does not rely on the context of sequences, while the second one takes into account that the features of layer-by-layer blocks might implicitly incorporate concepts of sequence modeling.
SVM Objective.
SVM aims to find an optimal hyperplane that separates classes by minimizing classification errors and maximizing the margin between support vectors.
CRF Objective.
CRF is used to capture sequence feature dependencies and make decisions based on neighboring element states in sequence labeling tasks, with the training objective of maximizing the conditional likelihood of the true label sequence given the input sequence.
5 Experiments
We now conduct experiments with AdaInfer on well-known LLMs across various tasks.
5.1 Evaluation Tasks
To evaluate the zero/few-shot learning capabilities of AdaInfer, we utilize three primary types of tasks.
Question Answering Tasks.
(1) MMLU Hendrycks et al. (2021) encompasses 57 tasks across humanities, social sciences, STEM, and more, requiring world knowledge and problem-solving capabilities. (2) CommonsenseQA Talmor et al. (2019) tests for commonsense knowledge through multiple-choice questions. (3) SQuAD Rajpurkar et al. (2016) serves as a reading comprehension benchmark, with questions based on Wikipedia articles and answers either segments of passage or marked as unanswerable.
Text Classification Tasks.
(1) SST-2 Socher et al. (2013) involves sentiment analysis of movie reviews with binary “positive” or “negative” labels. (2) AG News Zhang et al. (2015) classifies news headlines and article sentences into Business, Science/Technology, Sports, and World categories.
Rule Understanding Task.
GPT-3’s Brown et al. (2020) few-shot learning capability is tested with tasks requiring pattern recognition, using synthetic datasets from Todd et al. (2024); Hernandez et al. (2024) for tasks like Capitalize/Lowercase Letter, Choose Item/Category from List, and recognizing data pairs (e.g.,, Landmark-Country).
5.2 Experiment Settings
Table 1: LLMs statistics using AdaInfer.
Model Params Tokens Layer Num.
Meta/OPT 13B 0.18T 40
Meta/Llama 2 7B 2T 32
Meta/Llama 2 13B 2T 40
Meta/Llama 2 70B 2T 80
Table 2: Performance and efficiency in question answering tasks, with accuracy (%) denoted by ‘Acc’. Results include few-shot learning with sample sizes of 5, 10, 15, and 20, showcasing the average values.
Setting Model MMLU CommonsenseQA SQuAD Avg
[…]
Zero-shot OPT-13B 7.95 100 8.20 100 20.00 100 12.05 100
AdaInfer 8.67 97.55 2.80 97.55 23.00 97.55 11.49 97.55
Few-shot OPT-13B 23.60 100 21.45 100 26.12 100 23.72 100
AdaInfer 22.59 83.94 21.62 86.05 25.95 88.31 23.39 86.10
Zero-shot Llama2-13B 2.54 100 1.00 100 19.20 100 7.58 100
AdaInfer 2.48 98.14 0.70 98.37 25.90 85.34 9.69 93.95
Few-shot Llama2-13B 53.31 100 64.92 100 52.9 100 57.04 100
AdaInfer 52.44 93.55 62.48 89.10 48.35 80.66 54.42 87.77
Table 3: Performance and efficiency in classification and rule understanding, with accuracy (%) denoted by ‘Acc’. Results include few-shot learning with sample sizes of 5, 10, 15, and 20, showcasing the average values.
Setting Model Sentiment AG News Avg Rule Understanding
[…]
Zero-shot OPT-13B 0.00 100 0.10 100 0.05 100 3.38 100
AdaInfer 0.00 96.87 0.10 100 0.05 98.44 3.86 92.52
Few-shot OPT-13B 92.58 100 72.83 100 82.71 100 58.48 100
AdaInfer 92.97 80.28 72.83 100 82.90 90.14 52.83 89.74
Zero-shot Llama2-13B 0.00 100 0.10 100 0.05 100 2.32 100
AdaInfer 0.00 97.43 0.10 88.37 0.05 92.90 6.14 85.76
Few-shot Llama2-13B 95.90 100 77.53 100 86.72 100 69.36 100
AdaInfer 92.65 59.70 76.43 87.69 84.54 73.70 61.87 80.61
Large Language Models.
For AdaInfer’s backbone, we choose widely recognized LLMs, detailed in Table 1. These models vary in terms of the number of parameters, ranging from 7 billion to 70 billion, and the number of layers, ranging from 32 layers to 80 layers. Specifically, our selections encompass OPT Zhang et al. (2022) and the Llama 2 series Touvron et al. (2023). These models exhibit subtle differences in architectural design and training data volume.
In-context Learning Setting.
We evaluate our approach under zero-shot and few-shot scenarios, using sample sizes of 5, 10, 15, and 20. For zero-shot, the input is the test set’s $x_{q}$. For few-shot, training set examples are added to $x_{q}$. For in-context learning prompts, we use a default template: $\mathrm{Q}:\left\{x_{k}\right\}\backslash\mathrm{nA}:\left\{y_{k}\right\}\backslash\mathrm{n}\backslash\mathrm{n}$, concatenating random $x_{k}$ and $y_{k}$ samples from task-specific training sets.
Metrics.
For performance evaluation, we report the top-1 accuracy score on the test set following Todd et al. (2024). When tokenizing $y_{q}$ into multiple tokens, we treat the first token of $y_{q}$ as the target token. To evaluate computational efficiency, we determine the early exit layer index for each input instance, which can be translated into floating-point operations (FLOPs) ratios for comparison using the method described in Narayanan et al. (2021). The FLOPs ratio is calculated as:
$\frac{2l^{\prime}(6h+s)+V}{2l(6h+s)+V},$ (2)
where $l^{\prime}$ represents the stop layer index during inference in AdaInfer, $l$ is the total number of transformer layers, $h$ denotes the hidden size, $s$ is the sequence length, and $V$ stands for vocabulary size. Although AdaInfer converts hidden states to logits at each block through a classification layer, it only utilizes the last token’s hidden state even with longer sequences. Hence, this computation can be ignored (0.03% of the total FLOPs for transformer inference). Further computation details on the calculation process can be found in Appendix A. Statistical classifiers have significantly lower computational costs compared to LLM inference, as detailed in Appendix A, allowing us to overlook this aspect in our analysis.
5.3 Comparison with Baseline Methods
The main experimental results of AdaInfer are presented in Tables 2 and 3. These experiments were conducted in zero-shot and few-shot settings, showcasing the Top-1 accuracy and average FLOPs ratios (compared to the baseline). From a perspective of performance and computational efficiency, we can draw the following experimental conclusions.
Performance is Comparable with Minimal Loss.
Tables 2 and 3 show that across both zero-shot and few-shot settings, top-1 average accuracy remains within a narrow margin of $<$5% for all tasks and $<$1% for QA and text classification task groups when compared to baseline models. AdaInfer maintains LLMs capabilities and in-context learning abilities without modifying model parameters. This finding is promising, especially in light of our observation1 in Section 3.2, where we demonstrate the feasibility of implementing early exit strategies within LLM middle layers while preserving performance. For certain tasks, AdaInfer surpasses the last layer (baseline) in zero-shot or few-shot accuracy. This hints at a tendency for deep layers to potentially over-represent certain tasks, which could impede performance during LLM inference.
Reducing FLOPs Cost from 2% to 41%.
We convert the average and variance of early exit layers for each task to FLOPs ratios in Table 2 and Table 3. It can be observed that the FLOPs ratios vary for different types of tasks, ranging from $98\%$ to $59\%$. This variation is because AdaInfer assesses different early exit layer configurations for different task inputs. Even for the same task with different inputs, AdaInfer may recommend different early exit layer settings. For instance, in the sentiment analysis task, a $41\%$ reduction in computational cost can be achieved using Llama2-13B, while for the knowledge-based question answering MMLU and Commonsense question answering CommonSenseQA, the savings range from $2\%$ to $20\%$. This aligns with our observation2 outlined in Section 3.2, where we argue that at LLM inference scenario, Not all Layers are Necessary, and allocating fewer computational resources for “simple” samples can improve computational efficiency.
Table 4: Comparative analysis of GAP and CRF on performance and computational efficiency.
Task Setting AdaInfer w. Rule AdaInfer w. CRF
[…]
MMLU Zero-shot 5.35 90.84 4.77 97.40
Few-shot 47.09 84.10 52.72 97.15
CommonsenseQA Zero-shot 1.10 92.78 1.40 97.28
Few-shot 55.33 79.57 65.72 96.40
SQuAD Zero-shot 24.60 73.17 23.10 93.03
Few-shot 43.43 71.19 51.75 89.94
Sentiment Zero-shot 0.00 88.25 0.00 97.27
Few-shot 91.45 51.25 95.60 73.07
AG News Zero-shot 0.10 77.82 0.10 94.04
Few-shot 69.17 70.65 76.77 93.08
Rule Understanding Zero-shot 9.90 74.80 3.43 90.29
Few-shot 53.78 70.38 65.82 90.29
5.4 Evaluation on Different Exit Strategy
In the main experiments Table 2 and Table 3, we employ SVM as the classifier for AdaInfer. To explore the impact of different classification strategies, Table 4 compares the effects of implementing an early-exit strategy with a GAP threshold set at 0.8 (stopping computation when the current block’s GAP feature exceeds 0.8) against using CRF as a classifier. The results indicate that both GAP and CRF can reduce computational costs from 3% to 50% and maintain comparable LLM performance. Notably, in the zero-shot setting, GAP outperforms CRF, suggesting a relatively weak dependency between block features.
5.5 Evaluation across Scaling Law
In our main experiments, we employ 13B-sized Llama2 and OPT models. To explore the effects of AdaInfer on models of different sizes, we conduct experiments on the 7B and 70B versions of Llama2. The results for the 7B model, presented in Table 5, show that AdaInfer either maintains accuracy with minimal (<1%) loss or exceeds the baseline in certain tasks, and achieves a computational reduction ranging from 4% to 32%. However, in experiments with the 70B model, we observe that in a zero-shot setting, AdaInfer matches or slightly exceeds the baseline model while reducing computational costs by 10% to 50%. However, in the few-shot setting, despite similar reductions in computation, AdaInfer’s accuracy shows a 1% to 25% drop across different tasks compared to the baseline. This suggests that for larger models, such as the 70B or even larger scales, AdaInfer may need to more precisely identify and utilize features at different levels. Improving AdaInfer to adapt to these larger models is a direction for our future research. The results of all LLMs using different classifiers are summarized in Table 8 and Table 9 in the Appendix D and we have highlighted the best results for each task in the current setting.
Table 5: AdaInfer on Llama2-7B across tasks for performance and computational efficiency.
Task Setting Llama2-7B AdaInfer
[…]
MMLU Zero-shot 4.19 100 4.63 96.13
Few-shot 43.05 100 43.73 93.76
CommonsenseQA Zero-shot 5.30 100 4.80 95.26
Few-shot 53.50 100 53.00 90.46
SQuAD Zero-shot 20.40 100 23.80 89.98
Few-shot 48.08 100 45.82 87.06
Sentiment Zero-shot 0.00 100 0.00 96.37
Few-shot 95.20 100 95.30 68.05
AG News Zero-shot 0.10 100 0.10 91.36
Few-shot 79.65 100 79.72 94.51
Rule Understanding Zero-shot 5.47 100 5.32 91.55
Few-shot 66.80 100 66.92 88.41
5.6 Generalization Study
In Tables 2 and 3, we randomly select 6 to 9 training datasets from the entire pool of task training sets, which altogether contain 71 sub-datasets, to train the AdaInfer classifier. Furthermore, to assess the generalization performance of the statistical classifiers, we conduct the following tests.
  • Intra-Task Generalization. Evaluating the sentiment task using a classifier trained on the sentiment training dataset.
  • Inter-Task Generalization. Testing sentiment using a classifier trained on the knowledge question-answering task’s dataset.
  • Inter-Model Generalization. Assessing the sentiment task on Llama2-13B using a classifier trained on Llama2-7B.
The results are presented in Table 6. The SVM classifier exhibits satisfactory intra-task and inter-task generalization capabilities, consistent with the results presented in the main results. However, for the CRF classifier, training in an intra-task manner leads to premature termination of the LLM at very shallow layers, resulting in subpar performance. This could be attributed to insufficient feature selection, causing the CRF to overfit noise or local features in the training data. Additionally, due to variations in the logits distribution characteristics among different models, the inter-model classifier’s performance shows moderate accuracy. In conclusion, based on the results from Table 2 and Table 3 and Table 6, when using AdaInfer, we recommend utilizing SVM as the classifier.
Table 6: Generalization performance of statistic classifier on sentiment task on Llama2-7B (32 layers), Inter-Model refers to Llama2-13B (40 layers).
Classifier Generalization Acc Layers Variance FLOPs
SVM Intra-Task 94.90 18.15 0.45 60.58
CRF 0.00 0.00 0.00 0.00
SVM Inter-Task 95.50 19.20 4.40 63.80
CRF 94.90 20.20 4.55 66.87
SVM Inter-Model 90.70 20.60 3.70 54.55
CRF 87.75 19.20 2.75 51.09
5.7 Factor Study
In response to the features mentioned in Section 4.1, we conduct cross-validation. Given that the classifiers in the main results utilized basic features (i.e., “gap”, “top prob”), we explore the impact of features such as the cosine similarity between the current block and the previous block, which encompasses the attention values (attn), multi-layer perceptron (mlp), and hidden states. The results are presented in Table 7. The attention values have no discernible impact on the results, while other features like mlp and hidden states have an adverse effect. This result is consistent with the trend shown in Figure 3, indicating that logits can measure whether the model’s current forward progress is sufficient, while changes in other features may involve various factors.
Table 7: Comparative analysis of SVM performance with incremental feature addition in sentiment and MMLU/anatomy tasks.
Feature Sentiment MMLU
Base Features (gap, top prob) 94.90 41.13
+attn 94.90 41.13
+hidden state 67.53 41.13
+mlp 67.88 41.93
6 Conclusion
In this paper, we first give evidence of that Not all Layers are Necessary during Inference and provide statistical evidence to support this. Then, we present AdaInfer, a simple yet effective algorithm that determines the appropriate moment to cease inference based on the input instance, thus enhancing inference efficiency and adaptability without modifying the model’s parameters. Experiment results show that AdaInfer can reduce an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is compatible with other model acceleration techniques, potentially offering further improvements in inference efficiency. We argue that AdaInfer establishes a new paradigm for efficient inference besides effective existing methods.
Limitations
In this paper, we make the first attempt to discover that the logits of each block are critical for early-exit classifiers in LLMs, incorporating this insight as a key design choice in AdaInfer. However, since AdaInfer relies on a single forward pass, it has not yet been extended to sequential generative tasks, offering significant avenues for future research.
Ethics Statement
Our research aims to optimize large-scale model inference without modifying parameters, promising efficiency gains and reduced energy consumption. However, we must address potential misuse concerns, as enhanced inference capabilities may also enable malicious actors to exploit large neural language systems by injecting or amplifying logits as features, leading to undesirable behavior.
Acknowledgments
This work is supported by the National Science and Technology Major Project (2022ZD0116300) and the National Science Foundation of China (No. 62106249).
Appendix A Computation Cost.
Classifier Computation Cost.
We utilized the sklearn library for training SVM [^1] https://scikit-learn.org/stable/modules/svm.html and CRF [^2] https://sklearn-crfsuite.readthedocs.io/en/latest/, adhering to default configurations. For SVM and CRF training, we used the sklearn library with default settings. Given a training dataset with $N$ training examples, the time complexity for SVM training typically ranges from $O(N^{2}\times d)$ to $O(N^{3}\times d)$, where $d$ is the feature dimension. SVM prediction time complexity is $O(d)$ per single inference. For standard linear-chain CRF, the training time complexity is approximately $O(N\times S\times M)$, where $S$ is the average sequence length, $M$ is the label count. The prediction time complexity for CRF is $O(S\times M)$ per single inference. In contrast, the inference time complexity for large models like llama2 is $LSd(d+S)$ per single inference, where $d$ is the hidden size, $S$ is the sequence length, and $L$ represents the number of layers. Comparatively, the computational load of SVM and CRF is negligible when compared to large models.
Transformer Computation Cost.
Given a language model with $l$ transformer layers, hidden size $h$, sequence length $s$, vocabulary size $V$, and batch size $B$. Each transformer block needs $24Bsh^{2}+4Bs^{2}h$ FLOPs for the forward pass. The other main contributor to the FLOPs count is the classification layer in the language model head, which transforms features of dimension $h$ to the vocabulary dimension $V$. The required FLOPs for this operation is $2BshV$ in the forward pass. While AdaInfer does convert hidden states to logits at each block through classification layer, it only utilizes the hidden state from the last token for conversion, even when the sequence length is 2048 or longer. In the case of Llama2 7/13/70B, this computation accounts for only 0.000288, 0.000236, and 0.000152 of the total number of FLOPs for transformer inference. Similarly, for OPT 13B, it amounts to 0.000367. Consequently, the computational burden associated with this aspect can be disregarded. Summing these together, a transformer model with $l$ transformer layers, the total number of floating-point operations for inference is $4Bshl(6h+s)+2BshV$. Thus, the ratio of inference cost in FLOPs can be calculated through Equation 2.
Appendix B Details of Creating Training Data for Classifier
Considering a training input instance $x$ and its corresponding label $y$ from $D_{train}$. Once $x$ is processed through a decoder layer of LLM, we can extract a general feature vector $x^{d}$ ($d$ is the number of features). Additionally, we obtain the probability distribution $P$ over the vocabulary $V$ of the current layer’s hidden state after passing through the classification layer (as depicted in Section 3.1). This can be represented as: $P=\operatorname{softmax}(WH+b)$, where $H$ is the hidden state of the current layer, $W$ and $b$ are the weights and bias of the classification layer, respectively. Function $\operatorname{softmax}$ is applied to convert logits to probabilities. Let the highest-ranked token in this distribution be denoted as $\hat{y}=\text{argmax}(P)$, where $\text{argmax}(P)$ finds the token with the highest probability. If $\hat{y}$ matches the label $y$, the associated label $y_{c}$ for the feature vector $x_{d}$ is designated as positive; otherwise, it is labeled as negative. Thus, for an $L-$layer LLM, each input instance $x$ yields $L$ pairs of $<x^{d},y_{c}>$.
Appendix C More Observation of LLMs
Figure 4 depicts a visual analysis of features across the layers within each block of mainstream LLMs. It shows that the “gap” and “top prob” exhibit a gradual increase during the inference phase, reaching stability in the deeper layers. Additionally, the activation of “gap” and “top prob” varies across layers for different tasks. These observed trends align with the findings discussed in Section 4.1.
Appendix D Comprehensive Summary of Results
The results of all LLMs using different classifiers are summarized in Table 8 and 9. We have highlighted the best results for each task in the current setting. The experimental results indicate that (i) early exits are feasible for different tasks, (ii) the timing of early exits varies depending on the instance, and (iii) in both zero-shot and few-shot settings, accuracy is comparable with baseline models. It’s worth noting that for individual tasks, AdaInfer even outperforms the baseline in zero-shot or few-shot accuracy. This suggests that in inference scenarios, deep layers may tend to over-represent some tasks, potentially impairing performance.
Table 8: Performance and computational efficiency in question answering tasks, with accuracy (%) denoted by ‘acc’. Results include few-shot learning with sample sizes of 5, 10, 15, and 20, showcasing the average values.
Setting Model MMLU CommonsenseQA SQuAD Avg
[…]
Zero-shot OPT-13B 7.95 100 8.20 100 20.00 100 12.05 100
AdaInfer w. Rule 3.21 89.58 0.60 85.17 20.72 87.98 8.18 87.58
AdaInfer w. CRF 7.14 96.57 4.60 93.26 24.36 93.22 12.03 94.35
AdaInfer 8.67 97.55 2.80 97.55 23.00 97.55 11.49 97.55
Few-shot OPT-13B 23.60 100 21.45 100 26.12 100 23.72 100
AdaInfer w. Rule 20.99 79.54 20.72 80.00 24.20 82.93 21.97 80.82
AdaInfer w. CRF 24.44 97.43 21.18 97.55 25.98 97.11 24.81 97.37
AdaInfer 22.59 83.94 21.62 86.05 25.95 88.31 23.39 86.10
Zero-shot Llama2-7B 4.19 100 5.30 100 20.40 100 9.96 100
AdaInfer w. Rule 4.69 95.69 4.60 94.90 23.90 89.48 11.06 93.36
AdaInfer w. CRF 4.86 95.32 2.00 95.01 18.80 91.17 8.55 93.83
AdaInfer 4.63 96.13 4.80 95.26 23.80 89.98 11.08 93.79
Few-shot Llama-2-7B 43.05 100 53.50 100 48.08 100 48.21 100
AdaInfer w. Rule 44.03 93.69 52.83 90.23 45.68 86.72 47.51 90.21
AdaInfer w. CRF 41.38 94.23 53.6 91.61 43.62 88.10 46.20 91.31
AdaInfer 43.73 93.76 53.00 90.46 45.82 87.06 47.52 90.43
Zero-shot Llama2-13B 2.54 100 1.00 100 19.20 100 7.58 100
AdaInfer w. Rule 5.35 90.84 1.10 92.78 24.60 73.17 10.35 85.60
AdaInfer w.CRF 4.77 97.40 1.40 97.28 23.10 93.03 9.76 95.90
AdaInfer 2.48 98.14 0.70 98.37 25.90 85.34 9.69 93.95
Few-shot Llama-2-13B 53.31 100 64.92 100 52.9 100 57.04 100
AdaInfer w. Rule 47.09 84.10 55.33 79.57 43.43 71.19 48.62 78.29
AdaInfer w.CRF 52.72 97.15 65.72 96.40 51.75 89.94 56.73 94.50
AdaInfer 52.44 93.55 62.48 89.10 48.35 80.66 54.42 87.77
Table 9: Performance and computational efficiency in text classification and rule understanding tasks, with accuracy (%) denoted by ‘acc’. Results include few-shot learning with sample sizes of 5, 10, 15, and 20, showcasing the average values.
Setting Model Sentiment AG News Avg Rule Understanding
[…]
Zero-shot OPT-13B 0.00 100 0.10 100 0.05 100 3.38 100
AdaInfer w. Rule 0.00 90.61 0.10 92.03 0.05 91.32 3.64 87.55
AdaInfer w. CRF 0.00 97.55 0.10 97.55 0.05 97.55 4.11 97.55
AdaInfer 0.00 96.87 0.10 100 0.05 98.44 3.86 92.52
Few-shot OPT-13B 92.58 100 72.83 100 82.71 100 58.48 100
AdaInfer w. Rule 94.20 78.30 12.95 82.54 53.58 80.42 48.20 85.50
AdaInfer w. CRF 92.88 97.50 71.27 97.55 82.08 97.53 55.33 97.50
AdaInfer 92.97 80.28 72.83 100 82.90 90.14 52.83 89.74
Zero-shot Llama2-7B 0.00 100 0.10 100 0.05 100 5.47 100
AdaInfer w. Rule 0.00 96.08 0.10 91.05 0.05 93.57 5.41 91.20
AdaInfer w. CRF 0.00 96.07 0.10 92.20 0.05 94.14 3.62 92.08
AdaInfer 0.00 96.37 0.10 91.36 0.05 93.87 5.32 91.55
Few-shot Llama-2-7B 95.20 100 79.65 100 87.43 100 66.80 100
AdaInfer w. Rule 95.30 67.78 79.72 94.38 87.51 81.08 66.80 87.99
AdaInfer w. CRF 94.90 69.91 61.62 96.38 78.26 83.15 62.36 89.60
AdaInfer 95.30 68.05 79.72 94.51 87.51 81.28 66.92 88.41
Zero-shot Llama2-13B 0.00 100 0.10 100 0.05 100 2.32 100
AdaInfer w. Rule 0.00 88.25 0.10 77.82 0.05 83.04 9.9 74.80
AdaInfer w. CRF 0.00 97.27 0.10 94.04 0.05 95.66 3.43 90.29
AdaInfer 0.00 97.43 0.10 88.37 0.05 92.90 6.14 85.76
Few-shot Llama-2-13B 95.90 100 77.53 100 86.72 100 69.36 100
AdaInfer w. Rule 91.45 51.25 69.17 70.65 80.31 60.95 53.78 70.38
AdaInfer w. CRF 95.60 73.07 76.77 93.08 86.19 83.08 65.82 90.29
AdaInfer 92.65 59.70 76.43 87.69 84.54 73.70 61.87 80.61